# EVODiff: Entropy-aware Variance Optimized Diffusion Inference

### v2

@@ Abstract
Diffusion models (DMs) excel in image generation, but suffer from slow inference and the training-inference discrepancies. Although gradient-based solvers like DPM-Solver accelerate the denoising inference, they lack theoretical foundations in information transmission efficiency. In this work, we introduce an information-theoretic perspective on the inference processes of DMs, revealing that successful denoising fundamentally reduces conditional entropy in reverse transitions. This principle leads to our key insights into the inference processes: (1) data prediction parameterization outperforms its noise counterpart, and (2) optimizing conditional variance offers a reference-free way to minimize both transition and reconstruction errors. Based on these insights, we propose an entropy-aware variance optimized method for the generative process of DMs, called EVODiff, which systematically reduces uncertainty by optimizing conditional entropy during denoising. Extensive experiments on DMs validate our insights and demonstrate that our method significantly and consistently outperforms state-of-the-art (SOTA) gradient-based solvers. For example, compared to the DPM-Solver++, EVODiff reduces the reconstruction error by up to 45.5% (FID improves from 5.10 to 2.78) at 10 function evaluations (NFE) on CIFAR-10, cuts the NFE cost by 25% (from 20 to 15 NFE) for high-quality samples on ImageNet-256, and improves text-to-image generation while reducing artifacts. Code is available at https://github.com/ShiguiLi/EVODiff.
@@ 1 Introduction
  • Extensive experiments validate our insights and demonstrate significant improvements in inference. EVODiff reduces FID by 45.5% on CIFAR-10 (from 5.10 to 2.78 at 10 NFE) and 43.4% on LSUN-Bedrooms (from 13.97 to 7.91 at 5 NFE) over strong solvers like DPM-Solver++ and UniPC, while also mitigating visual artifacts in text-to-image generation.
@@ 5 Experiments
Model Method/NFE EMS-need? 5 8 10 12
@@
Model Method/NFE EMS-need? 5 10 15 20
@@ Conclusions
In this work, we propose EVODiff, a novel entropy-aware variance optimized method that addresses the inference process of diffusion models, significantly improving both efficiency and image generation quality. Specifically, our work first establishes a principled, information-theoretic foundation that explains why data-prediction parameterization outperforms its noise counterpart and demonstrates how optimizing conditional variance reduces transition and reconstruction errors without relying on reference trajectories. Building on these insights, EVODiff systematically reduces uncertainty in each denoising step, thereby accelerating convergence and significantly improving sample quality. Extensive experiments demonstrate EVODiff’s effectiveness across diverse settings with SOTA performance: on CIFAR-10, it outperforms the DPM-Solver++ baseline by 45.5% at 10 NFE (from 5.10 to 2.78 FID); on ImageNet-256, it reduces NFE cost by 25% (from 20 to 15 NFE) while maintaining high-fidelity generation; on LSUN-Bedrooms, it achieves up to 43.4% FID improvement over UniPC with 5.3% faster generation; and for text-to-image generation with Stable-Diffusion models, it produces superior visual quality while preserving semantic alignment. Our work achieves SOTA performance and introduces a reference-free variance-controlled inference framework for generative models, with the potential to address key challenges in advanced generative tasks.
@@
This work was supported in part by grants from the fundamental research program of Guangdong, China (2023A1515011281), the China Scholarship Council (202306150167), Guangdong Basic and Applied Basic Research Foundation (24202107190000687), Foshan Science and Technology Research Project (2220001018608).

### v3

@@ Abstract
Diffusion models (DMs) excel in image generation but suffer from slow inference and training-inference discrepancies. Although gradient-based solvers for DMs accelerate denoising inference, they often lack theoretical foundations in information transmission efficiency. In this work, we introduce an information-theoretic perspective on the inference processes of DMs, revealing that successful denoising fundamentally reduces conditional entropy in reverse transitions. This principle leads to our key insights into the inference processes: (1) data prediction parameterization outperforms its noise counterpart, and (2) optimizing conditional variance offers a reference-free way to minimize both transition and reconstruction errors. Based on these insights, we propose an entropy-aware variance optimized method for the generative process of DMs, called EVODiff, which systematically reduces uncertainty by optimizing conditional entropy during denoising. Extensive experiments on DMs validate our insights and demonstrate that our method significantly and consistently outperforms state-of-the-art (SOTA) gradient-based solvers. For example, compared to the DPM-Solver++, EVODiff reduces the reconstruction error by up to 45.5% (FID improves from 5.10 to 2.78) at 10 function evaluations (NFE) on CIFAR-10, cuts the NFE cost by 25% (from 20 to 15 NFE) for high-quality samples on ImageNet-256, and improves text-to-image generation while reducing artifacts. Code is available at https://github.com/ShiguiLi/EVODiff.
@@ 1 Introduction
  • Extensive experiments validate our insights and demonstrate significant improvements in inference. EVODiff reduces FID by 45.5% on CIFAR-10 (from 5.10 to 2.78 at 10 NFE) and 45.5% on LSUN-Bedrooms (from 13.97 to 7.91 at 5 NFE) over strong solvers like DPM-Solver++ and UniPC, while also mitigating visual artifacts in text-to-image generation.
@@ 5 Experiments
Model Method/NFE Entropy-aware? 5 8 10 12
@@
Model Method/NFE Reference-based? 5 10 15 20
@@ Conclusions
In this work, we propose EVODiff, a novel inference-time refinement method based on entropy-aware variance optimization. It significantly improves both efficiency and generative quality without relying on reference trajectories. Specifically, our work first establishes a principled, information-theoretic foundation that explains why data-prediction parameterization outperforms its noisy counterpart and demonstrates how optimizing conditional variance reduces transition and reconstruction errors without relying on reference trajectories. Building on these insights, EVODiff systematically reduces uncertainty in each denoising step, thereby accelerating convergence and significantly improving sample quality. Extensive experiments demonstrate EVODiff’s effectiveness across diverse settings with SOTA performance: on CIFAR-10, it outperforms the DPM-Solver++ baseline by 45.5% at 10 NFE (from 5.10 to 2.78 FID); on ImageNet-256, it reduces NFE cost by 25% (from 20 to 15 NFE) while maintaining high-fidelity generation; on LSUN-Bedrooms, it achieves up to 43.4% FID improvement over UniPC with 5.3% faster generation; and for text-to-image generation with Stable-Diffusion models, it produces superior visual quality while preserving semantic alignment. Our method achieves SOTA results and establishes a reference-free, variance-controlled inference framework, effectively addressing the trade-off between sampling efficiency and generative quality.
@@
This work was supported in part by grants from National Natural Science Foundation of China (52539005), the fundamental research program of Guangdong, China (2023A1515011281), the China Scholarship Council (202306150167), Guangdong Basic and Applied Basic Research Foundation (24202107190000687), Foshan Science and Technology Research Project(2220001018608).